OpenStudy (anonymous):

0.24(18) +0.04x=0.08(24+x)

OpenStudy (anonymous):

first reduce the bracket: 0.08(24+x)=0.08(24)+0.08x=1.92+0.08x next place like terms together: 0.24(18)-1.92 = 0.08x-0.04x 4.32-1.92=0.08x-0.04x 2.4=0.04x now divide the entire equation by 0.04 since you wish to find x: 2.4/0.04=0.04x/0.04(0.04/0.04=1;2.4/0.04=60 thus x = 60